OpenStudy (anonymous):

What is the slope of the line between (3, -4) and (-2, 1)?

OpenStudy (anonymous):

\[Slope = \frac{ 1-(-4) }{ -2-3 }\] \[=\frac{ 5 }{ -5 }\] = -1

OpenStudy (anonymous):

Alex is saving money because she would like to purchase the new video gaming system that everyone has been talking about. After two weeks of saving, she had $47. After 24 weeks of saving, she had $430. Approximately how many dollars per week did Alex save on average between weeks 2 and 24?

OpenStudy (anonymous):

$15.96 per week $17.41 per week $17.91 per week $19.88 per week

OpenStudy (anonymous):

Are you posting ur homework qns here?

OpenStudy (anonymous):

@J_Prince in the video gaming system question, you have two points you can make a line from the points and see where this line reaches the required value

OpenStudy (anonymous):

nahh a review

OpenStudy (anonymous):

The general formula is (final y-value - initial y-value)/(final x-value - initial x-value) x-value would be the independent variable, which is time in this case, and y-value would be the dependent variable.